OpenStudy (anonymous):

A pilot flew a 400-mile flight in 2.5 hours flying into the wind. Flying the same rate and with the same wind speed, the return trip took only 2 hours, with a tailwind. What was the speed of the wind?

OpenStudy (anonymous):

let the speed of plane=x m/hr speed of wind =y m/hr then (x-y)*(2.5)=400 x-y=400*(2/5)=160 ...(1) (x+y)*2=400 x+y=400/2=200 ...(2) subtract (1) from (2) and find y